# Co-Design of CNN Accelerators for TinyML using Approximate Matrix Decomposition

### 17 Apr 2026 v1

@@ II-A Approximate Weight Matrix Decomposition
Finally, Fig. 1c presents the decomposed version of a CNN convolution operation, assuming a convolutional kernel of size $N$ sliding over the input feature map. An input vector $In$ of size $N$ is also split into $S$ slices, such that $In=[In_{1}|In_{2}|...|In_{s}]$. A slice $In_{s}$ is then applied upon the decomposed matrices $F$ (corresponding to $W_{s}$) to produce $S$ partial sums. The convolution operation is then completed by accumulating the partial sums, as generated by the decomposed convolution, to produce $M$ outputs.
Figure 1: (a) CNN weight matrix transformation to $N$ and $M$ dimensions. (b) WMD of the transformed weight matrix to P decomposed F matrices with Po2 weights. (c) WMD multiplication with input and accumulation of the partial sums.
@@ III-A Processing Element (PE)
Moreover, we take advantage of some observed properties of the F{F} matrices: First, ${F}_{s,1}$ matrices always contain their non-zero elements in their first $S_{W}$ columns, which allows for omitting position encoding bits in the signals and multiplexers before the multiplier inputs. Second, in an effort to reduce indexing cost, we replace one of the non-zero elements $E$ with a fixed value of $1$ on the diagonal of the $F$ matrix (referred to as diagonal optimization hereafter). The activation is then passed directly to the adder tree, which allows for encoding E1E-1 elements per row. Then, all F-blocks corresponding to these matrices just need E1E-1 multiplier units per row, while each element of the input vector is hardwired to the adder tree of the corresponding row. Finally, for the purpose of lowering hardware resources, our design considers a set of Po2 with only negative exponents, which allows implementing right-shifts only.
@@ III-B Systolic Array (SA)
Our PEs are arranged in a 2-D grid in a pipelined manner, forming the core SA of our accelerator architecture, as shown in Fig. 2b. To map a convolution operation to our accelerator, we follow a process similar to Section II-A, but substituting $N$ for $C_{in}$. As can be seen in Fig. 2b, the input features reside in an on-chip input buffer, and then $S_{W}$ inputs flow from the top of the SA. Since each row corresponds to different output channels, the input features are shared along the y-axis. The partial sums are reduced along the x-axis via internal accumulators within each PE. They are then summed up with the outputs from the previous iteration and stored in an on-chip output buffer. The PEs in the x-axis are connected using $M$ adders. The parallelism of our SA (i.e., useful convolution MACs per cycle) can be determined based on the total number of PEs and the parallelism within each individual PE. At the PE level, this depends on the WMD parameters $S_{W}$ and $M$. Then the total number of outputs equivalent to MACs per cycle of our accelerator is $S_{W}\cdot M\cdot PE_{x}\cdot PE_{y}$, with $PE_{x}$ the number of columns and $PE_{y}$ the number of rows of the SA.
@@
To determine the 2D mapping $(PE_{x},PE_{y})$ on our SA, we use Algorithm 1. Starting from $PE_{x}=1$, the algorithm iteratively increments the length of the row, calculates $PE_{y}$ based on the total number of LUTs of the FPGA, and computes the latency $Lat$, based on the hardware cost of a single PE unit $R_{PE_{unit}}$ and the cost of a PE row $R_{PE_{x}}$, which can be derived from the F-blocks cost $R_{F_{gen}}$ and $R_{F_{0}}$ as described in Section IV-B. The process continues until the cost of one row exceeds the available LUTs. Once the algorithm terminates, the best latency $Lat_{best}$ is returned, ensuring that for a given CNN, the assignment best suited to its characteristics is chosen. This ensures that the best PE mapping is chosen for a given CNN based on the $C_{out},C_{in}$ dimensions.
@@ V-A Experimental Setup
We evaluate our framework on the set of CNNs available in MLPerfTiny [3], a widely used TinyML benchmark that comprises representative CNN workloads customized to the stringent TinyML resource constraints. As CNN benchmarks, it includes ResNet, MobileNetV1, and DS-CNN, trained on CIFAR-10, VWW, and Speech Commands, respectively. For the framework evaluation, we exclusively use the pre-trained models of all three CNNs from MLPerfTiny, without additional retraining or finetuning. TFLite is used to evaluate the accuracy of both our decomposed and the baseline CNNs. We evaluate the solutions found with our framework against traditional state-of-the-art SA-based accelerators [17] that use low precision MAC units (i.e., weights from 4 to 8 bits) across all MLPerfTiny CNNs. The mapping of each CNN to these SAs is obtained using Algorithm 1. Each row and column of the SA can be supplied with weights and activations in a single clock cycle through appropriately instantiated BRAM-based input and output buffers. As our target platform, we use the Artix-7 XC7A100TCSG324-1 FPGA. A fully parametrized SA using our WMD PEs as well as the SAs using MAC units as PEs are implemented in SystemVerilog. Vivado 2023.1 is used for simulation and synthesis to obtain FPGA resource utilization in terms of LUTs, flip-flops (FFs), and BRAMs, as well as to estimate power. For our resource model, we synthesize its basic blocks and extract their resource utilization in terms of LUTs (see Section IV-B1). The final latency is calculated for each CNN on a layer-wise basis using Vivado simulation, from which toggle rates are also obtained to serve as inputs for the power consumption analysis. Functional verification uses uniformly random inputs, decomposed weights, and $8$-bit activations.
Our NSGA-II is implemented using pymoo [4], with an initial population of $250$ chromosomes, running for up to $20$ generations. We set the crossover probability and degree of mutation $\eta_{c}$ to $0.9$ and $15$, respectively, and the mutation $\eta_{m}$ to $5$. For the genetic exploration we reserve $10\%$ of the test set for the exploration itself and $90\%$ of the test set to evaluate the final accuracy of our designs. For our largest benchmark (i.e., MobileNet), the size of the explored design space—as defined by the WMD parameters $P_{h},P_{s}$—reaches for $P_{s}=54$, $2^{12}\cdot 54=221\,184$ possible solutions, highlighting the need for employing genetic-based exploration within our framework.
@@ V-B Evaluation
Number of FFs (overall utilization) 32301 (25%) 27578 (22%) 25900 (20%) 21233 (17%) 22927 (18%) 19033 (15%)
@@
Number of FFs (overall utilization) 32301 (25%) 27578 (22%) 25427 (20%) 21233 (17%) 22618 (18%) 13674 (11%)
@@
Number of FFs (overall utilization) 32301 (25%) 27578 (22%) 25545 (20%) 21233 (17%) 23170 (18%) 18390 (17.5%)
@@ V-C Comparison Against State-of-the-Art Low-Precision MAC-based Systolic Array
We compare our approach against PTQ, a state-of-the-practice technique for improving hardware efficiency [10]. Using the same baseline SA architecture and adjusting the precision of the MAC unit, we progressively quantize the weights of the evaluated CNNs to lower bit-widths while keeping activations fixed at $8$ bits. Figure 5 shows the achievable peak throughput of the MAC-based SA at various bit-widths, normalized with respect to the accelerators obtained with our framework. Overall, we observe that PTQ at bit-widths lower than $5$ bits incurs substantial accuracy loss, well over the $2\%$ accuracy drop threshold. At the same time, our approach consistently outperforms PTQ within the $2\%$ accuracy drop threshold showcasing the effectiveness of our Po2-based framework in providing throughput improvements at a minimal loss in terms of accuracy. Table II, Table III, and Table IV contain the PTQ obtained top-1 accuracies, hardware resources and performance metrics for the different weight resolution systolic arrays. The best values for each row are highlighted for an easier comparison. The 4-bit SA, although providing advantages in many hardware performance metrics, causes a top-1 accuracy drop of at least 6% for the analyzed CNNs, which is why this configuration is excluded in the following comparisons. It can be observed that the DS-CNN obtained configuration outperforms the SAs across most of the hardware performance metrics. In general, the slightly higher estimated power consumption of the obtained accelerators is compensated by the higher peak throughput, which leads to lower latencies and higher energy efficiency in terms of GOPS/W. Figure 6 shows that the accelerator solutions found with our framework lie in the Pareto front of all evaluated CNNs, demonstrating optimal trade-offs between latency, accuracy, and energy efficiency.
@@ V-D Comparison Against State-of-the-Art Post-Training Po2 FPGA Implementation
Finally, we compare our framework methodology to the state-of-the-art Po2 acceleration technique, ShiftCNN, proposed by [11] [11]. We re-implement ShiftCNN targeting the same FPGA as in our work and by precisely following the design directives detailed in [11]. This approach approximates a normalized weight by adding $N$ numbers from a set of negative powers of 2, each selected by a $B$-bit index. Then, the hardware architecture includes a precomputed tensor storing right-shifted input activations by different amounts, for which just one is selected using encoded $N\cdot B$-bit weights as inputs to multiplexers. A parameter $C$ is also considered in the architecture to define the number of weight and activation pairs that the accelerator can receive as inputs per cycle, which at the same time defines the number of multiplexers instantiated. While not explicitly described, $N\cdot C$ multiplexers are needed to keep the multiplication result in a single cycle. All multiplexers’ outputs are accumulated with an adder tree.

### 23 Apr 2026 v2

@@ II-A Approximate Weight Matrix Decomposition
Finally, Fig. 1c presents the decomposed version of a CNN convolution operation, assuming a convolutional kernel of size NN sliding over the input feature map. An input vector InIn of size NN is also split into $S$ slices, such that $In=[In_{1}|In_{2}|...|In_{s}]$. A slice $In_{s}$ is then applied upon the decomposed matrices $F$ (corresponding to $W_{s}$) to produce $S$ partial sums. The convolution operation is then completed by accumulating the partial sums, as generated by the decomposed convolution, to produce $M$ outputs.
Figure 1: (a) CNN weight matrix transformation to NN and $M$ dimensions. (b) WMD of the transformed weight matrix to P decomposed F matrices with Po2 weights. (c) WMD multiplication with input and accumulation of the partial sums.
@@ III-A Processing Element (PE)
Moreover, we take advantage of some observed properties of the F{F} matrices: First, ${F}_{s,1}$ matrices always contain their non-zero elements in their first $S_{W}$ columns, which allows for omitting position encoding bits in the signals and multiplexers before the multiplier inputs. Second, in an effort to reduce indexing cost, we replace one of the non-zero elements $E$ with a fixed value of 11 on the diagonal of the $F$ matrix (referred to as diagonal optimization hereafter). The activation is then passed directly to the adder tree, which allows for encoding E1E-1 elements per row. Then, all F-blocks corresponding to these matrices just need E1E-1 multiplier units per row, while each element of the input vector is hardwired to the adder tree of the corresponding row. Finally, for the purpose of lowering hardware resources, our design considers a set of Po2 with only negative exponents, which allows implementing right-shifts only.
@@ III-B Systolic Array (SA)
Our PEs are arranged in a 2-D grid in a pipelined manner, forming the core SA of our accelerator architecture, as shown in Fig. 2b. To map a convolution operation to our accelerator, we follow a process similar to Section II-A, but substituting NN for $C_{in}$. As can be seen in Fig. 2b, the input features reside in an on-chip input buffer, and then $S_{W}$ inputs flow from the top of the SA. Since each row corresponds to different output channels, the input features are shared along the y-axis. The partial sums are reduced along the x-axis via internal accumulators within each PE. They are then summed up with the outputs from the previous iteration and stored in an on-chip output buffer. The PEs in the x-axis are connected using $M$ adders. The parallelism of our SA (i.e., useful convolution MACs per cycle) can be determined based on the total number of PEs and the parallelism within each individual PE. At the PE level, this depends on the WMD parameters $S_{W}$ and $M$. Then the total number of outputs equivalent to MACs per cycle of our accelerator is $S_{W}\cdot M\cdot PE_{x}\cdot PE_{y}$, with $PE_{x}$ the number of columns and $PE_{y}$ the number of rows of the SA.
@@
To determine the 2D mapping $(PE_{x},PE_{y})$ on our SA, we use Algorithm 1. Starting from $PE_{x}=1$, the algorithm iteratively increments the length of the row, calculates $PE_{y}$ based on the total number of LUTs of the FPGA, and computes the latency LatLat, based on the hardware cost of a single PE unit $R_{PE_{unit}}$ and the cost of a PE row $R_{PE_{x}}$, which can be derived from the F-blocks cost $R_{F_{gen}}$ and $R_{F_{0}}$ as described in Section IV-B. The process continues until the cost of one row exceeds the available LUTs. Once the algorithm terminates, the best latency $Lat_{best}$ is returned, ensuring that for a given CNN, the assignment best suited to its characteristics is chosen. This ensures that the best PE mapping is chosen for a given CNN based on the $C_{out},C_{in}$ dimensions.
@@ V-A Experimental Setup
We evaluate our framework on the set of CNNs available in MLPerfTiny [3], a widely used TinyML benchmark that comprises representative CNN workloads customized to the stringent TinyML resource constraints. As CNN benchmarks, it includes ResNet, MobileNetV1, and DS-CNN, trained on CIFAR-10, VWW, and Speech Commands, respectively. For the framework evaluation, we exclusively use the pre-trained models of all three CNNs from MLPerfTiny, without additional retraining or finetuning. TFLite is used to evaluate the accuracy of both our decomposed and the baseline CNNs. We evaluate the solutions found with our framework against traditional state-of-the-art SA-based accelerators [17] that use low precision MAC units (i.e., weights from 4 to 8 bits) across all MLPerfTiny CNNs. The mapping of each CNN to these SAs is obtained using Algorithm 1. Each row and column of the SA can be supplied with weights and activations in a single clock cycle through appropriately instantiated BRAM-based input and output buffers. As our target platform, we use the Artix-7 XC7A100TCSG324-1 FPGA. A fully parametrized SA using our WMD PEs as well as the SAs using MAC units as PEs are implemented in SystemVerilog. Vivado 2023.1 is used for simulation and synthesis to obtain FPGA resource utilization in terms of LUTs, flip-flops ($F$s), and BRAMs, as well as to estimate power. For our resource model, we synthesize its basic blocks and extract their resource utilization in terms of LUTs (see Section IV-B1). The final latency is calculated for each CNN on a layer-wise basis using Vivado simulation, from which toggle rates are also obtained to serve as inputs for the power consumption analysis. Functional verification uses uniformly random inputs, decomposed weights, and 88-bit activations.
Our NSGA-II is implemented using pymoo [4], with an initial population of $250$ chromosomes, running for up to 2020 generations. We set the crossover probability and degree of mutation $\eta_{c}$ to $0.9$ and $15$, respectively, and the mutation $\eta_{m}$ to 55. For the genetic exploration we reserve $10\%$ of the test set for the exploration itself and $90\%$ of the test set to evaluate the final accuracy of our designs. For our largest benchmark (i.e., MobileNet), the size of the explored design space—as defined by the WMD parameters $P_{h},P_{s}$—reaches for $P_{s}=54$, $2^{12}\cdot 54=221\,184$ possible solutions, highlighting the need for employing genetic-based exploration within our framework.
@@ V-B Evaluation
Number of $F$s (overall utilization) 32301 (25%) 27578 (22%) 25900 (20%) 21233 (17%) 22927 (18%) 19033 (15%)
@@
Number of $F$s (overall utilization) 32301 (25%) 27578 (22%) 25427 (20%) 21233 (17%) 22618 (18%) 13674 (11%)
@@
Number of $F$s (overall utilization) 32301 (25%) 27578 (22%) 25545 (20%) 21233 (17%) 23170 (18%) 18390 (17.5%)
@@ V-C Comparison Against State-of-the-Art Low-Precision MAC-based Systolic Array
We compare our approach against PTQ, a state-of-the-practice technique for improving hardware efficiency [10]. Using the same baseline SA architecture and adjusting the precision of the MAC unit, we progressively quantize the weights of the evaluated CNNs to lower bit-widths while keeping activations fixed at 88 bits. Figure 5 shows the achievable peak throughput of the MAC-based SA at various bit-widths, normalized with respect to the accelerators obtained with our framework. Overall, we observe that PTQ at bit-widths lower than 55 bits incurs substantial accuracy loss, well over the $2\%$ accuracy drop threshold. At the same time, our approach consistently outperforms PTQ within the $2\%$ accuracy drop threshold showcasing the effectiveness of our Po2-based framework in providing throughput improvements at a minimal loss in terms of accuracy. Table II, Table III, and Table IV contain the PTQ obtained top-1 accuracies, hardware resources and performance metrics for the different weight resolution systolic arrays. The best values for each row are highlighted for an easier comparison. The 4-bit SA, although providing advantages in many hardware performance metrics, causes a top-1 accuracy drop of at least 6% for the analyzed CNNs, which is why this configuration is excluded in the following comparisons. It can be observed that the DS-CNN obtained configuration outperforms the SAs across most of the hardware performance metrics. In general, the slightly higher estimated power consumption of the obtained accelerators is compensated by the higher peak throughput, which leads to lower latencies and higher energy efficiency in terms of GOPS/W. Figure 6 shows that the accelerator solutions found with our framework lie in the Pareto front of all evaluated CNNs, demonstrating optimal trade-offs between latency, accuracy, and energy efficiency.
@@ V-D Comparison Against State-of-the-Art Post-Training Po2 FPGA Implementation
Finally, we compare our framework methodology to the state-of-the-art Po2 acceleration technique, ShiftCNN, proposed by [11] [11]. We re-implement ShiftCNN targeting the same FPGA as in our work and by precisely following the design directives detailed in [11]. This approach approximates a normalized weight by adding NN numbers from a set of negative powers of 2, each selected by a $B$-bit index. Then, the hardware architecture includes a precomputed tensor storing right-shifted input activations by different amounts, for which just one is selected using encoded $N\cdot B$-bit weights as inputs to multiplexers. A parameter $C$ is also considered in the architecture to define the number of weight and activation pairs that the accelerator can receive as inputs per cycle, which at the same time defines the number of multiplexers instantiated. While not explicitly described, $N\cdot C$ multiplexers are needed to keep the multiplication result in a single cycle. All multiplexers’ outputs are accumulated with an adder tree.